OpenStudy (anonymous):

Which equation supports the given conjecture? The sum of the squares of two consecutive integers is an odd integer. A. (2 + 3)2 B. 32 + 52 = 34 C. 4 + 52 = 29 D. 52 + 62 = 61

OpenStudy (anonymous):

@redtornado

OpenStudy (emmamink):

sorry girl, no can help..@confluxepic

OpenStudy (emmamink):

@confluxepic

OpenStudy (confluxepic):

None of them are correct.

OpenStudy (confluxepic):

@Rapunzel101402 @EmmaMink @TheSmartOne @~

OpenStudy (emmamink):

How do u know?

TheSmartOne (thesmartone):

Well people if they copy and paste their questions and don't spend any time to look at it, we can't help them.

TheSmartOne (thesmartone):

The person could have atleast fixed it by putting in the squares.

TheSmartOne (thesmartone):

The options should have been \(\Huge A.(2+3)^2\) \(\Huge B. 3^2+5^2=34\) \(\Huge C. 4+5^2=29\) \(\Huge D. 5^2+6^2=61\) If these are the correct answer choices then the correct answer is D.

TheSmartOne (thesmartone):

Because it says the sum of the squares has to equal an odd integer. So the sum of the squares of 5 and 6 does equal to an odd integer.

TheSmartOne (thesmartone):

@rapunzel101402 The CoC says http://prntscr.com/5a6u4k Next time make sure your question is clear so that others can help you without getting all confused. http://openstudy.com/code-of-conduct And since you are new, please read the CoC ^^ and feel free to read this. http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8